OpenStudy (anonymous):

Consider the following problem: Florence has 8 nieces and she wants to divide her collection of dolls among them evenly. If each niece receives thirty-two dolls, how many dolls were in Florence's collection? If d represents the number of dolls in Florence's collection, which of the following equations can be used to model the problem? 8 times d equals thirty-two thirty-two times d equals 8 thirty-two divided by 8 equals d d divided by 8 equals thirty-two

OpenStudy (anonymous):

I think its D) d divided by 8 equals thirty-two, although, B) seems correct as well

OpenStudy (anonymous):

I think its D as well

OpenStudy (anonymous):

D

OpenStudy (anonymous):

It's D, yeah If you attach "units" to to the numbers it can make it a little more cleasr. d = Total number of dolls 32 = dolls / niece 8 = Nieces \[ \frac{8 \ \ nieces}{1} \frac{32 \ dolls}{niece} = d \ \ total \ \ dolls \] 8 * 32 = d

OpenStudy (anonymous):

Okay, @AllTehMaffs is really smart so his answer the best ^ :D

OpenStudy (anonymous):

y'alls had already figured it out ^_^

OpenStudy (anonymous):

Great answer @AllTehMaffs :D